# ExBody2: Advanced Expressive Humanoid Whole-Body Control

### v1

@@ 3.2.2 Student Policy Training
To train the student, we adopt the strategy used in DAgger [59], we roll out the student policy $\pi$ in the simulation environment to generate training data. For each visited state, the teacher policy $\hat{\pi}$ computes the oracle action as the supervision signal. We proceed to refine the policy $\pi$ by iteratively minimizing the loss $l$ on the accumulated data. The training of $\hat{\pi}$ continues through successive rollouts until the loss $l$ reaches convergence. A critical aspect of training the student policy is preserving a sufficiently long sequence of historical observations. Detailed results and further analysis are elaborated in Section 4.
@@
Table 1: Rewards Specification for ExBody2.
3.2.3 Decomposed Tracking Strategy
Motion tracking comprises two objectives: tracking DoF (joint) positions and keypoint (body keypoint) positions. Keypoint tracking usually plays a crucial role in tracking motions, as joint DoF errors can propagate to the whole body, while keypoint tracking is directly applied to the body. Existing work like H2O, OmniH2O [24, 23] learns to follow the trajectory of global keypoints. However, this global tracking strategy usually results in suboptimal or failed tracking behavior, as global keypoints may drift through time, resulting in cumulative errors that eventually hinder learning. To address this, we map global keypoints to the robot’s current coordinate frame, and instead utilize velocity-based global tracking. The coordination of velocity and motion allows tracking completion with maximal expressiveness, even if slight positional deviations arise. Moreover, to further enhance the robot’s capabilities in following challenging keypoint motions, we allow a small global drift of keypoints during training stage and periodically correct them to the robot’s current coordinate frame. During deployment, we strictly employ local keypoint tracking with velocity-decomposed control.
3.3 Continual Motion Synthesis
The existing motions are usually short in length, which limits humanoid robots from performing interesting and extended behaviors continuously. To overcome this, we train a Conditional Variational Autoencoder (CVAE) [64, 79] to synthesize future movements that can be directly employed by the whole-body tracking policy, allowing the robot to execute complex and expressive movements seamlessly during deployment.
As shown in Figure 2 (c), the CVAE takes on past $M-$step motions $m_{t-M:t}$, and synthesizes future $H$-step motions $m_{t+1:t+H+1}$ autoregressively. $m_{t}$ includes the current DoF positions, root pose, velocity, and angular velocity. Specifically, we utilize a transformer-based CVAE architecture, similar to the architecture proposed in  [79]. We first separately tokenize DoF positions and root information through two separate MLPs, and get tokenized motion information. We compute the posterior $q_{\phi}(m_{t-M:t},m_{t+1:t+H+1})$ through a transformer encoder network, using the features from the [$CLS$] token. For notational simplicity and consistency with standard CVAE, we denote the condition $c=m_{t-M:t}$ to be the history context, and $x=m_{t+1:t+H+1}$ to be the prediction target. During training time, the latent variable $z$ is sampled from $\mathcal{N}(\mu(c,x),\sigma(c,x)^{2})$. A transformer decoder will take $z$ with $c=m_{t-M:t}$, and positional embeddings to predict $H-$step future motions $\hat{x}$. The CVAE loss consists of the reconstruction loss $l_{recon}=\|x-\hat{x}\|^{2}$ and the KL divergence loss:
$l_{KL}=\frac{1}{2}\sum_{i=1}^{d}\left(\sigma_{i}^{2}(x,c)+\mu_{i}^{2}(x,c)-1-\log(\sigma_{i}^{2}(x,c))\right)$
where $d$ is the dimensionality of the latent space. To improve the smoothness of the predicted motion, we also apply a smoothness loss:
$l_{smooth}=\|\hat{m}_{t+1}-m_{t}\|^{2}+\sum_{i=t+1}^{t+H}\|\hat{m}_{i+1}-\hat{m}_{i}\|^{2}$
The total loss for cVAE training is: $l_{motion}=l_{recon}+\alpha l_{KL}+\beta l_{smooth}$ where we set $\alpha=\beta=0.5$ in our experiments. During inference, $z$ is set to the mean of the prior (i.e. zero). We adopt a similar temporal ensemble strategy as in  Zhao et al. [79] to get predicted future motions. Further details regarding the encoder and decoder architecture of the CVAE are in the supplementary materials.
4 Experiments
Method Sim2Real $E_{\text{vel}}\downarrow$ $E_{\text{mpkpe}}\downarrow$ $E_{\text{mpkpe}}^{\text{upper}}\downarrow$ $E_{\text{mpkpe}}^{\text{lower}}\downarrow$ $E_{\text{mpjpe}}\downarrow$ $E_{\text{mpjpe}}^{\text{upper}}\downarrow$ $E_{\text{mpjpe}}^{\text{lower}}\downarrow$
Robot G1
@@ 4 Experiments
Exbody 0.3083 0.1134 0.0924 0.1343 0.1660 0.0766 0.2891
OmniH2O* 0.2429 0.0959 0.0878 0.1041 0.1103 0.0913 0.1364
Exbody2 0.1891 0.0854 0.0781 0.0928 0.0938 0.0784 0.1149
Robot H1
@@
Exbody 0.5964 0.1941 0.1948 0.1916 0.3251 0.2784 0.3775
OmniH2O* 0.4254 0.1540 0.1563 0.1509 0.2184 0.2135 0.2240
Exbody2 0.3016 0.1290 0.1317 0.1247 0.1763 0.1722 0.1809
Table 2: Simulation motion imitation evaluation of Exbody2 and baselines on dataset $\mathcal{D}_{250}$ for robots G1 and H1. Statistically significant results are highlighted in bold across 5 random seeds.
4.1 Experimental Setup
We conduct our experiments in IsaacGym [46] simulator across two robot platforms (Unitree G1 and H1).
Baselines. To examine the effectiveness of varying tracking methods, motion control strategies, and training techniques, we evaluate four baselines using a high-quality, curated dataset of 250 motion samples. This dataset is carefully selected for optimal training suitability, as will be shown in subsequent experiments.
@@ 4.1 Experimental Setup
  • Exbody [9]: This method only tracks the upper body movements from the human data, while tracking the root motion of the lower body without explicitly following step patterns and focusing on partial body tracking.
[…]
  • OmniH2O*: Our reimplementation of the OmniH2O [23], using global keypoints tracking and the same observation space as described in the original paper. During deployment, we adapt OmniH2O to our local tracking evaluation for fair comparison.
  • ExBody2: Our method utilizes local keypoint tracking. It incorporates various training techniques to enhance the overall motion fidelity and sim-to-real transfer.
@@
Metrics
(a) Eval. on $\mathcal{D}_{50}$
$\mathcal{D}_{50}$ 0.1375 0.0627 0.0571 0.0682 0.0753 0.0626 0.0928
$\mathcal{D}_{250}$ 0.1454 0.0669 0.0600 0.0738 0.0870 0.0689 0.1119
$\mathcal{D}_{500}$ 0.1593 0.0740 0.0660 0.0820 0.1051 0.0865 0.1307
(b) Eval. on $\mathcal{D}_{\text{CMU}}$
$\mathcal{D}_{50}$ 0.2621 0.1076 0.1071 0.1078 0.1343 0.1321 0.1373
$\mathcal{D}_{250}$ 0.1901 0.1048 0.1009 0.1084 0.1119 0.0998 0.1286
$\mathcal{D}_{500}$ 0.1966 0.1083 0.1044 0.1124 0.1252 0.1106 0.1452
$\mathcal{D}_{\text{CMU}}$ 0.1788 0.1106 0.1047 0.1171 0.1308 0.1130 0.1552
(c) Eval. on $\mathcal{D}_{\text{ACCAD}}$
$\mathcal{D}_{50}$ 0.4226 0.1277 0.1210 0.1330 0.1720 0.1618 0.1861
$\mathcal{D}_{250}$ 0.3533 0.1234 0.1141 0.1315 0.1421 0.1223 0.1692
$\mathcal{D}_{500}$ 0.3581 0.1280 0.1198 0.1355 0.1586 0.1400 0.1842
Table 3: Dataset Ablation Study: Evaluation on $\mathcal{D}_{50}$, $\mathcal{D}_{\text{CMU}}$, and $\mathcal{D}_{\text{ACCAD}}$ datasets with models trained on various datasets. Statistically significant results are highlighted in bold across 5 random seeds.
4.2 Main Results
Table 2 and Figure 3 present the results of different algorithms on two robot platforms, Unitree G1 and H1, where ExBody2 outperforms all other methods in all the tracking metrics. Our approach maintains consistent lower tracking error across the entire time frame with minimal fluctuations, indicating a stable and precise tracking performance. To be more precise, we evaluate the following metrics:
@@ 4.2 Main Results
    Upper Body Tracking Accuracy: Our algorithm performs slightly lower than Exbody in upper body tracking accuracy. However, Exbody sacrifices lower body tracking to maintain balance, whereas our algorithm can perform well with balanced tracking accuracy for both upper body and lower body.
@@
    Lower Body Tracking Accuracy: Our algorithm demonstrates higher accuracy in lower body tracking compared to other algorithmes for both G1 and H1.
@@
    Velocity Tracking Accuracy: Our algorithm also excels in velocity tracking accuracy, outperforming other methods. This improvement is attributed to the teacher-student training paradigm, where the distillation from privileged information to historical observation results in a student policy with better velocity-tracking.
Overall, our algorithm achieves significant improvements in full-body tracking accuracy for both upper and lower body, and velocity tracking accuracy compared to the baseline algorithms, demonstrating stable and effective tracking performance in dynamic environments.
@@
(a) History Length Ablation
Exbody2-History25 (Ours) 0.1891 0.0854 0.0781 0.0928 0.0938 0.0784 0.1149
Exbody2-History0 0.3101 0.0896 0.0812 0.0980 0.0972 0.0805 0.1201
Exbody2-History10 0.1934 0.0860 0.0792 0.0927 0.0935 0.0777 0.1152
Exbody2-History50 0.1944 0.0871 0.0797 0.0945 0.0962 0.0791 0.1198
Exbody2-History100 0.2119 0.0854 0.0778 0.0929 0.0960 0.0787 0.1198
(b) Reset Global Tracking Ablation
Exbody2(Ours) 0.1891 0.0854 0.0781 0.0928 0.0938 0.0784 0.1149
Exbody2-w/o-Reset 0.2545 0.0871 0.0807 0.0935 0.0919 0.0752 0.1149
(c) DAgger Ablation
Exbody2(Ours) 0.1891 0.0854 0.0781 0.0928 0.0938 0.0784 0.1149
Exbody2-w/o-DAgger 0.3115 0.0925 0.0843 0.1006 0.1152 0.0996 0.1366
Table 4: Self Ablation Study: Evaluation of different configurations of our method on dataset $\mathcal{D}_{250}$. The table is divided into three parts: (a) History Length Ablation, (b) Reset Global Tracking Ablation, and (c) DAgger Ablation. The best results in each column are highlighted.
4.3 Ablation on Datasets
We manually design four datasets of varying sizes, where the largest being the complete CMU AMASS dataset, containing in total 1,919 motions. The remaining datasets, with sizes 50, 250, and 500, are subsets of the CMU dataset, each constructed with different levels of action diversity:
  • 50-action dataset ($D_{50}$): This subset includes only fundamental static actions, such as standing and walking. Both upper and lower limb actions lack diversity, primarily covering basic and easily feasible motions.
  • 250-action dataset ($D_{250}$): In addition to basic actions, this subset introduces more diverse upper limb actions and a variety of lower limb actions, such as running and simple dance movements. These actions are feasible for robots, albeit moderately challenging.
  • 500-action dataset ($D_{500}$): This subset includes increasingly complex lower limb actions, such as single-leg jumps and intricate dance moves, rendering the dataset somewhat noisier with less feasible motions.
  • Full CMU dataset ($D_{CMU}$): The complete CMU dataset comprises 1,919 actions, including extreme movements like push-ups, rolling on the ground, and somersaults. This dataset is unfiltered and introduces significant noise.
We train ExBody2 on the above four datasets, and conduct three sets of evaluations on the G1 robot, shown in Table 3:
@@ 4.3 Ablation on Datasets
    Evaluation on $D_{50}$: Policies trained on $D_{50}$ achieve the highest tracking accuracy for in-distribution actions, as reflected in metrics across all categories. This suggests that additional data does not necessarily benefit in-distribution tasks. While the policy trained on $D_{250}$ performs similarly to $D_{50}$, policies trained on $D_{500}$ and $D_{CMU}$ exhibit substantial drops in tracking accuracy.
@@
    Evaluation on $D_{CMU}$: Policies trained on $D_{250}$ achieve the best performance on $D_{CMU}$, surpassing those trained on both $D_{500}$ and the full $D_{CMU}$ dataset. Due to the limited diversity of the $D_{50}$ dataset, especially in upper limb movements, the $D_{50}$-trained policy struggles to maintain high accuracy for out-of-distribution actions. Unexpectedly, the $D_{250}$-trained policy generalizes better than those trained on $D_{500}$ and $D_{CMU}$. This result underscores that noisy datasets degrade policy performance, as the policy may expend unnecessary effort on tracking infeasible actions, lowering the accuracy of feasible actions.
  3. 3.
    Evaluation on $D_{ACCAD}$: This experiment further emphasizes the importance of clean datasets. Here, the ACCAD dataset ($D_{ACCAD}$) consists of actions that are entirely not in the training data. Policy trained on $D_{250}$ outperforms the others, achieving the best tracking accuracy. Additionally, the $D_{250}$, $D_{500}$, and $D_{CMU}$-trained policies perform relatively well in velocity tracking. However, the $D_{50}$-trained policy suffers from substantial tracking errors, suggesting the limitations of a small, simple dataset in handling unseen data.
In conclusion, the $D_{250}$ dataset yields the best performance, providing a balance between diversity and feasibility. It includes a rich variety of upper limb actions and moderately challenging lower limb actions that remain feasible for robots. This dataset avoids the drawbacks of overly simplistic datasets with insufficient diversity and noisy datasets that degrade tracking performance.
4.4 Ablation on Policy Training
Finally, we ablate the design choices in ExBody2, including the effectiveness of the teacher-student setup, the optimal history length to train the policy, and the mechanism of using delayed keypoints reset, as shown in Table 4. The experiments are conducted on the Unitree G1 robot.
We first examine the impact of teacher-student training. As shown in Table 4 (c), the tracking accuracy drops significantly without teacher-student training. This is largely due to the absence of privileged velocity guidance, making it difficult for the single-stage RL policy to learn velocity directly from history. Consequently, it could not keep up with the desired movements, leading to worse tracking precision.
We then test student policies trained with different history lengths in Table 4 (a). When no extra history is used, the policy struggles to learn effectively. Among the non-zero history lengths, most policies perform similarly while the history length of 25 yields the best results, which is used by us in the main experiments. Longer history lengths increase the difficulty of fitting the privileged information, ultimately reducing tracking performance.
Lastly, we evaluate the impact of adding small drifting by using a delayed global keypoint reset strategy in Table 4 (b). We allow the keypoints to move in the global frame for short periods before resetting to the robot’s local frame. This effectively serves as a data augmentation strategy, encouraging the robot to track keypoints with more significant drift. This design notably improves velocity tracking, enabling the robot to use absolute position adjustments to compensate for discrepancies when tracking velocities.
5 Conclusion
In this work, we develop ExBody2, a novel algorithm that achieves stable and expressive whole-body motion for humanoid robots by utilizing a curated dataset and a decoupled whole-body tracking framework. Our approach separates key body tracking from velocity control, enabling precise and stable motion replication. While we have identified the characteristics of datasets best suited for humanoid robot training, an automated method to select high-quality datasets remains an open challenge. Addressing this limitation will be a focus of our future work.

### v2

@@ III-B2 Student Policy Training
To train the student, we adopt the strategy used in DAgger [50], we roll out the student policy $\pi$ in the simulation environment to generate training data. For each visited state, the teacher policy $\hat{\pi}$ computes the oracle action as the supervision signal. We proceed to refine the policy $\pi$ by iteratively minimizing the loss $l$ on the accumulated data. The training of $\hat{\pi}$ continues through successive rollouts until the loss $l$ reaches convergence. A critical aspect of training the student policy is preserving a sufficiently long sequence of historical observations. Detailed results and further analysis are elaborated in Section IV.
@@
TABLE I: Rewards Specification for Exbody2.
III-B3 Motion-velocity Decoupled Control Strategy
Motion tracking comprises two objectives: tracking DoF (joint) positions and keypoint (body keypoint) positions. Keypoint tracking usually plays a crucial role in tracking motions during training stage, as joint DoF errors can propagate to the whole body, while keypoint tracking is directly applied to the body. Existing work like H2O, OmniH2O [18, 17] learns to follow the trajectory of global keypoints. However, this global tracking strategy usually results in suboptimal or failed tracking behavior, as global keypoints may drift through time, resulting in cumulative errors that eventually hinder learning. To address this, we map global keypoints to the robot’s current coordinate frame, and instead utilize velocity-based global tracking. The coordination of velocity and motion allows tracking completion with maximal expressiveness, even if slight positional deviations arise. Moreover, to further enhance the robot’s capabilities in following challenging keypoint motions, we allow a small global drift of keypoints during training stage and periodically correct them to the robot’s current coordinate frame. During deployment, we strictly employ local keypoint tracking with motion-velocity decoupled control.
IV Experiments
In this section, we present experiments to evaluate the effectiveness of Exbody2. We first introduce the experimental setup and the baselines, followed by a detailed analysis addressing the following key research questions:
Q1. (Section IV-B) Does Exbody2 generalist policy achieve higher tracking accuracy in both simulation and real-world deployment compared to prior methods?
Q2. (Section IV-C) What selection criteria lead to the optimal subset of a human motion dataset for learning a better generalist policy?
Q3. (Section IV-D) Does finetuning a specialist policy for specific motion groups further improve tracking performance?
Exbody 0.4700 0.1339 0.1249 0.1428 0.2020 0.1343 0.2952
@@ IV Experiments
OmniH2O* 0.3725 0.1253 0.1266 0.1240 0.1681 0.1564 0.1843
Exbody2-w/o-Filter 0.2787 0.1133 0.1087 0.1182 0.1355 0.1192 0.1579
Exbody2(Ours) 0.2930 0.1000 0.0960 0.1040 0.1079 0.0953 0.1253
TABLE II: Comparisons with baselines on dataset $\mathcal{D}_{CMU}$ for Unitree G1. We loop each motion trajectory 10 times in simulation and compute the per-step average error for each policy over all repetitions. The lowest error in each column for each group is highlighted in bold.
Method $E_{\text{mpjpe}}\downarrow$ $E_{\text{mpjpe}}^{\text{upper}}\downarrow$ $E_{\text{mpjpe}}^{\text{lower}}\downarrow$
Exbody 0.2178 0.1223 0.3239
@@
OmniH2O* 0.1396 0.1273 0.1533
Exbody2-w/o-Filter 0.1361 0.1254 0.1481
Exbody2(Ours) 0.1074 0.1092 0.1054
TABLE III: Comparisons with baselines on selected motions for Unitree G1 in real world.
IV-A Experimental Setup
Baselines. To examine the effectiveness of varying tracking methods, motion control strategies, and training techniques, we evaluate four baselines using a classic CMU dataset [1], which includes a wide variety of action types.
  • Exbody [4]: This method utilizes one-stage RL training pipeline, and only tracks the upper body movements from the human data, while tracking the root motion of the lower body without explicitly following step patterns and focusing on partial body tracking.
@@ IV-A Experimental Setup
  • OmniH2O*: Our reproduction of the OmniH2O [17], using global keypoints tracking and the same observation space as described in the original paper. During deployment, we adapt OmniH2O to our local tracking evaluation for fair comparison.
  • Exbody2: Our method utilizes local keypoint tracking and employs data curation to refine the training set. It incorporates various training techniques to enhance overall motion fidelity and improve sim-to-real transfer.
@@
IV-B Generalist Policy Performance
[…]
The only trade-off observed is a slight increase in velocity tracking error compared to the unfiltered version. We attribute this to the broader velocity patterns found in the full dataset, which facilitate more diverse dynamic behaviors but also introduce additional noise. Overall, the modest velocity trade-off is outweighed by significant enhancements in stability and precision. These trends are observed under simulation conditions, and we further analyze real-world performance in the following section.
In real-world experiments, we selected a representative subset from the CMU dataset, encompassing a diverse range of actions—such as expressive standing postures, walking at various speeds and patterns, squatting, and dancing. Using the robot’s onboard motor readings, we computed the joint position error, as reported in Table III. The experimental outcomes closely mirror the simulation results, demonstrating that our algorithm achieves higher tracking accuracy compared to other baselines for both upper and lower body segments. Notably, the integration of automated data curation has substantially enhanced performance. This improvement is critical in real-world environments, where unanticipated disturbances are more prevalent, underscoring the importance of maintaining robust and consistent behavior to achieve high-precision tracking throughout the motion sequences.
Overall, our generalist policy achieves significant improvements in full-body tracking accuracy for both upper and lower body, and velocity tracking accuracy compared to the baseline algorithms both in simulation and the real world, demonstrating stable and effective tracking performance in dynamic environments.
IV-C Impact of Automatic Data Curation
Figure 4: Impact of dataset filtering thresholds on policy tracking errors. The figure shows the tracking error trends across different dataset filtering thresholds. Policies trained on datasets with filtering thresholds that balance diversity and stability (e.g., $\pi_{\tau=0.150}$) achieve the lowest tracking errors. The base policy exhibits suboptimal performance due to unfiltered data, while overly restrictive thresholds (e.g., $\pi_{\tau=0.075}$) and overly lenient thresholds (e.g., $\pi_{\tau=0.175}$) show reduced effectiveness. We compute the error metric $e(s)=\alpha\,E_{\text{key}}(s)+\beta\,E_{\text{dof}}(s)$ with $\alpha=0.1,\beta=0.9$, assigning heavier weight to the joint-angle term.
To experimentally examine the selection criteria for constructing an optimal human motion dataset that enhances generalist policy learning, we conducted the following experiment, reconstructing the full pipeline of our automated data curation method and evaluating its effect on policy performance.
@@ IV-C Impact of Automatic Data Curation
    Base Policy Training: We first trained a base policy on the unfiltered $\mathcal{D}_{CMU}$ dataset, which encompasses a broad range of human motions, including both stable and highly dynamic sequences, as well as infeasible movements that exceed the robot’s physical limitations.
@@
    Dataset Filtering: The base policy’s tracking performance was evaluated on $\mathcal{D}_{CMU}$, and each motion sequence was assigned a tracking error score. Based on these scores, we sorted the sequences and applied filtering thresholds to create progressively refined datasets: $\mathcal{D}_{\tau=0.075}$, $\mathcal{D}_{\tau=0.1}$, $\mathcal{D}_{\tau=0.125}$, $\mathcal{D}_{\tau=0.15}$, $\mathcal{D}_{\tau=0.175}$. A lower threshold retains only motions with the lowest tracking errors, ensuring high stability but reducing diversity. Conversely, a higher threshold allows more challenging and diverse motions but introduces tracking instability.
    These thresholds were determined based on the error distribution of the base policy, as detailed in the appendix Section 4.B. The distribution-informed selection ensures that the filtering process is not arbitrary but rather grounded in empirical data. This selection strategy is not limited to a specific dataset; rather, it generalizes to other datasets with similar policy architectures and training conditions. When applied to new datasets, it effectively filters motions while maintaining a balance between feasibility and diversity.
@@
    Filtered Policy Resume: For each filtered dataset $\mathcal{D}_{\tau}$, we resume training from the base policy $\pi_{0}$ to obtain a new policy $\pi_{\tau}$. Concretely, thresholds $\tau\in\{0.075,0.1,0.125,0.15,0.175\}$ each yield a distinct subset $\mathcal{D}_{\tau}$, and thus a correspondingly refined policy $\pi_{\tau}$. This resumption process leverages the base model’s learned prior, enabling faster adaptation to each subset’s motion characteristics and improving overall training efficiency.
  4. 4.
    Evaluation: All the resulting policies were evaluated on the full $\mathcal{D}_{CMU}$ dataset, measuring tracking performance across multiple metrics. The results are visualized in Figure 4.
The results confirm that dataset selection significantly influences both generalization and tracking accuracy. A well-chosen dataset, striking a balance between stability and diversity, produces more robust and adaptable policies, in line with our feasibility-diversity principle.
  • Low Thresholds (e.g., $\pi_{\tau=0.075}$): Policies trained on heavily filtered datasets exhibited poor generalization, as they primarily learned from static and simple motions. While these datasets ensured stability, the lack of diversity limited the policy’s ability to handle more complex behaviors.
  • High Thresholds (e.g., $\pi_{\tau=0.175},\pi_{0}$): Policies trained on datasets with high thresholds struggled due to the inclusion of highly dynamic and unstable motions. The increased data variability introduced noise during training, leading to inconsistent policy behavior and reduced tracking accuracy.
  • Moderate Thresholds (e.g., $\pi_{\tau=0.15}$): Policies trained on datasets with intermediate thresholds achieved the best trade-off between feasibility and diversity. The dataset $\mathcal{D}_{\tau=0.15}$ retained sufficient variability to improve generalization while excluding excessively difficult or unstable motions, leading to the lowest overall tracking error.
These findings validate the importance of carefully balancing feasibility and diversity in dataset selection. The dataset $\mathcal{D}_{\tau=0.15}$ and its corresponding policy $\pi_{\tau=0.15}$ emerged as the optimal choice, ensuring both strong generalization and stable tracking performance. This highlights the necessity of structured dataset curation in developing robust robotic policies, particularly for real-world deployment where adaptability and reliability are crucial.
IV-D Specialist Policy finetuning
@@
(b) $\mathcal{D}_{easy}$
Specialist 0.0828 0.0561 0.0564 0.0558 0.0772 0.0647 0.0944
Scratch 0.0853 0.0608 0.0623 0.0592 0.0843 0.0711 0.1024
Generalist 0.0986 0.0699 0.0708 0.0690 0.1041 0.0882 0.1259
(a) $\mathcal{D}_{Moderate}$
Specialist 0.0991 0.0571 0.0582 0.0559 0.0760 0.0636 0.0930
Scratch 0.1188 0.0676 0.0688 0.0663 0.0924 0.0794 0.1103
Generalist 0.1217 0.0741 0.0727 0.0755 0.1092 0.0914 0.1337
(c) $\mathcal{D}_{Hard}$
Specialist 0.1712 0.0827 0.0829 0.0826 0.1047 0.0911 0.1234
Scratch 0.1631 0.0886 0.0898 0.0873 0.1188 0.1067 0.1354
Generalist 0.1452 0.0890 0.0867 0.0912 0.1181 0.1011 0.1414
(d) $\mathcal{D}_{ACCAD}$
Specialist 0.4021 0.1149 0.1079 0.1215 0.1402 0.1290 0.1557
Scratch 0.4153 0.1246 0.1154 0.1332 0.1609 0.1490 0.1771
Generalist 0.3361 0.1268 0.1156 0.1391 0.1716 0.1532 0.1967
TABLE IV: Comparison of three training strategies (Generalist, Specialist, Scratch) across four dataset groups ($\mathcal{D}_{easy}$, $\mathcal{D}_{moderate}$, $\mathcal{D}_{hard}$, and $\mathcal{D}_{ACCAD}$).
Figure 5: A sequence of a robot performing the Cha-Cha dance. From top to bottom: the reference motion represented by an avatar, our algorithm’s performance in the simulation, and its performance on a real robot. The bottom three rows show the per-frame errors: whole-body joint DoF error, upper-body joint DoF error, and lower-body DoF error, with the blue curve representing Exbody2-Specialist policy finetuned on $\mathcal{D}_{dancing}$ , orange for Exbody2-Scratch policy training from scratch on $\mathcal{D}_{dancing}$, green for our Exbody2-Generalist policy trained on filtered $\mathcal{D}_{CMU}$.
We evaluate the effectiveness of the pretrain-finetune paradigm in enhancing policy performance. To this end, we compare three distinct training strategies:
  • Generalist: A generalist policy trained on the automatically filtered dataset ($\pi_{\tau=0.15}$), designed to provide broad motion coverage and strong generalization across diverse tasks.
  • Specialist: A specialist policy obtained by fine-tuning the pretrained generalist policy on task-specific datasets, enabling higher precision for specialized motions.
  • Scratch: A policy trained from scratch on the same task-specific datasets. To ensure fairness, we match its total training iterations to the sum of pretraining plus fine-tuning iterations used in the Specialist approach.
To assess the robustness and adaptability of these policies, we conduct experiments across four manually curated datasets:
  • $\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{moderate}},\mathcal{D}_{\text{hard}}$: A series of datasets with increasing difficulty levels, categorized based on motion dynamics. Lower-difficulty datasets mainly contain static or low-movement motions, while higher-difficulty datasets include more dynamic and high-momentum movements. This progression allows us to assess how well the policies generalize to increasingly complex motions.
  • $\mathcal{D}_{\text{ACCAD}}$: An out-of-distribution (OOD) dataset used to evaluate the generalization capability of the learned policies on previously unseen motion patterns.
The quantitative evaluation results are summarized in Table IV. The key findings from these experiments are discussed below.
@@
    Performance on $\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{moderate}},\mathcal{D}_{\text{hard}}$: The finetuned policy achieves the best performance across all difficulty levels. The more challenging the dataset, the greater the advantage of finetuning over training from scratch, demonstrating the importance of leveraging a pretrained policy as a foundation for specialized tasks. For highly dynamic motions, the generalist policy exhibits slightly better velocity tracking due to its broader exposure to diverse movements. However, the specialist policy consistently achieves higher overall precision.
@@
    Performance on $\mathcal{D}_{\text{ACCAD}}$: The finetuned policy significantly outperforms both the pretrained and scratch policies on the OOD dataset, highlighting its superior generalizability and adaptability to unseen scenarios. This result further confirms that the pretrained generalist policy provides a strong foundation, while finetuning enhances task-specific adaptation.
To better illustrate the effectiveness of the specialist policy, we select the Cha-Cha dance as a case study, as shown in Figure 5. Cha-Cha is one of the representative motions in the dance group and involves dynamic lower-body movements combined with expressive upper-body gestures. By comparing policies trained from scratch, the generalist policy, and the specialist policy fine-tuned on the dancing dataset, we observe that the specialist policy achieves significantly lower tracking errors across all key metrics. This result further confirms the advantage of task-specific finetuning in capturing fine-grained motion details while maintaining stability.
In conclusion, the pretrain-finetune paradigm proves to be an effective strategy for achieving robust and adaptable policies. The pretrained generalist policy ($\pi_{\tau=0.15}$) provides a strong starting point, while finetuning allows specialization for specific tasks, resulting in superior performance across diverse datasets. This approach demonstrates significant benefits, particularly in challenging and OOD scenarios, and highlights the importance of combining generalist capabilities with task-specific specialization.
V Conclusion
This paper introduces Advanced Expressive Whole-Body Control (Exbody2), a novel framework for humanoid whole-body control that achieves superior tracking accuracy, stability, and adaptability through automated dataset filtering, a generalist-specialist training pipeline, and a decoupled keypoint-velocity tracking strategy. Experiments show that Exbody2 outperforms prior methods by balancing feasibility and diversity in dataset selection, enabling more robust whole-body tracking, and improving generalization across diverse motion tasks. The specialist finetuning further refines performance for high-precision tasks, demonstrating the effectiveness of a structured pretrain-finetune paradigm. These contributions push the boundaries of humanoid motion control, paving the way for more expressive and stable real-world deployments.
VI Limitations
Although our generalist policy already provides broad coverage of diverse motions, it does not capture the fine-grained precision of each specialist policy within a single framework. One limitation of our approach is the inability to seamlessly recombine policies that have been fine-tuned for specific motion groups. This constraint reduces the flexibility of switching between different policies within a single tracking session, as each specialist policy remains focused on a particular subset of motions. Consequently, transitions between different motion types may not be handled as smoothly or efficiently as desired.
In principle, unifying the broad coverage of the generalist policy with the high-accuracy tracking of individual specialists could provide the best of both worlds. Addressing this limitation by incorporating a dynamic policy-integration mechanism—where specialized policies can be adaptively blended or switched in real time—could significantly improve overall tracking precision, adaptability, and robustness, especially when dealing with complex, multi-modal motion sequences.